In a sample of 1000 workers, if 200 of them are graduates, then what is the sample proportion of graduate workers? In addition, if 650 of them are married, what is the proportion of married workers?
1
Expert's answer
2020-05-14T16:49:05-0400

1)

n=1000, N=200

sample proportion of graduate workers:

μp=p=N/n=0.2

σp="\sqrt{(p(1-p))/n}= 0.0126"

2)n=1000, N=650

sample proportion of married workers:

μp=p=N/n=0.65

σp="\sqrt{(p(1-p))/n}=0.015"